OpenStudy (anonymous):

Elimination Method a+8b=16 2+2b=-2 got (4,-3) Can someone tell me if that is correct?

jimthompson5910 (jim_thompson5910):

is the 2nd equation equation really 2+2b = -2?

OpenStudy (anonymous):

That was a mistake it is a+2b=-2

jimthompson5910 (jim_thompson5910):

No, (4,-3) is not the correct answer

OpenStudy (anonymous):

can you explain it to me step by step. I am not getting the correct answer

jimthompson5910 (jim_thompson5910):

Notice we have 'a' in the first and second equations. So we can multiply both sides of the second equation with -1 and add to get rid of these terms a+2b=-2 -1(a+2b)=-1(-2) -a-2b = 2 So we now have a+8b=16 -a-2b = 2 Now add the two equations a+8b=16 -a-2b = 2 ---------- 0a+6b = 18 So 6b = 18

jimthompson5910 (jim_thompson5910):

Solve for b. Then use this to find 'a'

OpenStudy (anonymous):

so i got b=3 a=-8 Correct?

jimthompson5910 (jim_thompson5910):

yes you got it

OpenStudy (anonymous):

In this equation will this be the correct form (3,-8)?

jimthompson5910 (jim_thompson5910):

no, put it in alphabetical order

jimthompson5910 (jim_thompson5910):

so 'a' comes first, then b

jimthompson5910 (jim_thompson5910):

that's why they call it an *ordered* pair

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

yw